Simultaneous equation with logarithims:
Solve for x and y the following:
Log(2x-5y) = 0; 2Logy = Log(x+1)

Answer by Alan3354(69443)

Solve for x and y the following:
Log(2x-5y) = 0;
--> 2x - 5y = 1
---------------
2Logy = Log(x+1)
log(y^2) = log(x+1)
y^2 = x+1
x = y^2 - 1
-------
2x - 5y = 1
Sub for x
2(y^2-1) - 5y = 1
2y^2 - 5y -3 = 0
(2y + 1)*(y - 3) = 0
y = 3, x = 8
==============
y = -1/2, x = -3/4